You have sold short 50 shares of company LOL at $82. Your initial margin is 50% and your maintenance margin is 30%. A few weeks later the company paid $2 dividends and the stock price hit $89. What is your percentage margin?
Answer #1

Initially money in your account=50*82*50%=$2050

When stock hit $89, total loss=(89-82)*50=$350

Money in your account=2050-350=$1700

So, Percentage margin=1700/(89*50)=0.3820 or 38.2%